An investment pays 3.4% compounded four times a year.
a. What is the annual growth factor?
Round to four decimal places.


b. What is the annual growth rate?
Round to two decimal places.

Answer by stanbon(75887)

An investment pays 3.4% compounded four times a year.
a. What is the annual growth factor?
Round to four decimal places.
(r/n)^4 = (1+(0.034/4))^4 = 1.0344
--------------------------------------

b. What is the annual growth rate?
Round to two decimal places.
0.0344 = 3.44%
-------------
Cheers,
Stan H.